# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v2

@@ Abstract
Although magnetic field line chaos, cantori, and turnstiles underlie the physics of tokamak disruptions, runaway electron damage, stellarator non-resonant divertors, and the most important electromagnetic correction to what are called electrostatic micro-instabilities, these concepts are not well known. These concepts will be defined and applications that illustrate their importance will be discussed.
@@ 1 Introduction
Positions in space are fundamentally defined by Cartesian coordinates, $\vec{x}=x\hat{x}+y\hat{y}+z\hat{z}$. Any three quantities $(\psi,\theta,\varphi)$ can be used as coordinates by defining the position function $\vec{x}(\psi,\theta,\varphi)$ as long as the Jacobian
$\mathcal{J}\equiv\left(\frac{\partial\vec{x}}{\partial\psi}\times\frac{\partial\vec{x}}{\partial\theta}\right)\cdot\frac{\partial\vec{x}}{\partial\varphi}$ (1)
is neither zero nor infinity in the spatial region in which they are to be used. Defining the position function just means giving $x$, $y$, and $z$ as functions of $(\psi,\theta,\varphi)$. The appendix to (Boozer:2004) explains how arbitrary coordinate systems, such as $\vec{x}(\psi,\theta,\varphi)$, are used in calculations.
@@
(Boozer:1983) showed the topological properties of magnetic field lines at each instant in time are given by $\psi_{p}(\psi_{t},\theta,\varphi,t)$. The position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ is needed to plot the field lines in ordinary space with $\partial\vec{x}/\partial t$ giving the velocity of $(\psi_{t},\theta,\varphi)$ through space, which is the field line velocity when field line topology is not changing, which implies $\psi_{p}$ can be taken to be independent of time.
When a particular line $\vec{x}_{0}(\ell)$ remains within a bounded region of space, it has three possibilities as $\ell\rightarrow\infty$: It can close on itself after a distance $\ell_{0}$. It can cover a surface coming arbitrarily close to every point on a surface without ever going through the same point twice. It can cover a volume coming arbitrarily close to every point in a volume of space without ever going through the same point twice. All three types of trajectories are illustrated by the Standard Map, Figure 2. The iteration number of the map represents the distance $\ell$. The periodicity of both coordinates of the Standard Map makes it non-trivial to translate them into the $(\psi_{t},\theta)$ coordinates of the magnetic field line Hamiltonian $\psi_{p}$. (Boozer-Punjabi:2018) in their Appendix B have obtained an explicit formula for a $\psi_{p}(\psi_{t},\theta,\varphi)$ that represents the qualitative features of stellarator equilibria including the non-resonant divertor that (Strumberger:1992) found naturally arises at the edge optimized stellarators. Non-resonant divertors are discussed in Section 5.2 as way to build intuitive understanding of the subtleties of cantori and turnstiles.
@@
The behavior of the magnetic field lines that are infinitesimally separated from an arbitrary line $\vec{x}_{0}(\ell)$ can be studied without approximation by a particularly simple Hamiltonian (Boozer:2012). The study of neighboring lines, which is reviewed in Appendix A, showed: (1) It is possible for field lines infinitesimally separated from a given line $\vec{x}_{0}(\ell)$, of any of the three types, to have separations that increase or decrease exponentially with $\ell$. There must be an equal flux of neighboring lines that diverge from and approach the line $\vec{x}_{0}(\ell)$ due to $\vec{\nabla}\cdot\vec{B}=0$. When lines that lie on a circle of infinitesimal radius at $\ell=0$ are followed a distance $\ell$, the circle will deform into an ellipse. The field strength times the square of the semi-major axis of the ellipse increases as $e^{\sigma_{max}(\ell)}$, and the field strength times the square semi-minor axis decreases as $e^{-\sigma_{max}(\ell)}$. (2) When the magnetic field depends on only two spatial coordinates, a single field line cannot cover a volume, as in the red area of Figure 2. Three spatial coordinates are required. Two-coordinate magnetic fields, as in axisymmetry, can have X-points (more correctly X-lines) which are specific field lines from which neighboring lines exponentially approach and separate. However, the volume in which lines have at least a certain exponentiation is itself exponentially small. (3) Curl-free magnetic fields can exhibit field lines with an exponential separation throughout a volume. An ideal magnetic evolution does not conserve the exponentiation $\sigma_{max}(\ell)$. In other words, the exponentiation can be a property of the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ with $\psi_{p}$ an unchanging function of its canonical coordinates. The current density required for an ideal evolution to produce an exponentiation $\sigma_{max}$ is proportional to $\sigma_{max}$, Appendix A.3. As will be discussed in Section 6, the time required for a large scale reconnection to occur is determined by a timescale set by the ideal evolution times $\sigma_{r}$. The required exponentiation $\sigma_{r}$ for large scale reconnection is the ratio of timescale for non-ideal diffusive effects, such as resistivity, to a timescale determined by the ideal evolution. As discussed in Section 6.2, the time scale for reconnection is determined by an ideal timescale times the natural logarithm of $\sigma_{r}$.
@@
The tendency to construct two dimensional models has led to the prejudice that exponential separation of magnetic field lines throughout a volume is peculiar. Actually any magnetic field that depends on all three spatial coordinates tends to be chaotic, which is illustrated by the design of a curl-free magnetic fields for a stellarators. Obtaining nested magnetic surfaces, as required for a stellarator, rather than having the field lines fill a volume requires careful optimization. Even a small perturbation to an optimized curl-free stellarator can destroy the surfaces. Nevertheless, axisymmetric tokamaks are even more sensitive to external field perturbations due to their amplification by coupling to marginally stable current-driven kinks (Park:2010).
@@
(Boozer:2004) showed the time derivative of the poloidal flux is the loop voltage,
@@
At the magnetic axis, Equation (2) is a trivial consequence of Faraday’s Law plus Stokes’ Theorem with the loop voltage $V_{\ell}$ the integral of the electric field around the axis. The proof of this equation for an arbitrary magnetic surface, which is given in the Appendix to (Boozer:2004), is non-trivial. On an arbitrary magnetic surface, the loop voltage is an integral along a field line at an instant of time,
@@
As proven in the paragraph below Equation (17), when $\vec{\nabla}V_{\ell}=0$ the topology of the magnetic field lines cannot change, and their evolution is considered ideal. This is in contrast to ideal magnetohydrodynamics, which requires a zero resistivity $\eta$, and that implies $V_{\ell}=0$. As pointed out by (Boozer:2025:Constraints) and discussed in Section 7.1, the distinction between $\vec{\nabla}V_{\ell}=0$ and $V_{\ell}=0$ is important when devising strategies for avoiding disruptions in tokamaks. The condition $\vec{\nabla}V_{\ell}=0$ is consistent with changing the magnetic flux at the axis of a toroidal plasma without changing the magnetic field in the plasma while the condition $V_{\ell}=0$ is not.
@@
Non-resonant perturbations can be taken to change the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ while the poloidal flux can be held fixed using a canonical transformation or what is mathematically equivalent an electric potential. This is shown in the appendix to (Boozer:2004).
Section 4 discusses the subtle properties of chaotic magnetic field lines. These properties need to be understood before undertaking a study of cantori and turnstiles, which serve as boundaries on region of chaotic magnetic field lines.
(Chirikov:1960) gave the approximate condition for the creation of large chaotic volumes: the overlap of the islands from different resonant rational surfaces.
As discussed in Section 5, the actual formation of chaotic regions is more complicated than just the overlap of island chains. (MacKay:1984) showed that chaotic regions are formed by cantori and the turnstiles that penetrate them.
@@
The highly collimated flux tubes of turnstiles are responsible for the severe damage that can be caused by runaway electrons following the lines in these tubes to localized positions on the chamber walls (Boozer-Punjabi:2016). This collimation also responsible for the formation of a natural non-resonant divertor at the edge of a stellarator (Strumberger:1992) and (Boozer-Punjabi:2018). Section 5.2 on non-resonant divertors provides an intuitive introduction to cantori and turntiles through an application to a practical problem.
If the plasma evolution were perfectly ideal, a resonant perturbation would affect only the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$. But, a resonant ideal perturbation forces the magnetic field to have a delta-function current density at the resonant rational surface $\iota=n/m$. This current causes distortions to the neighboring surfaces producing magnetic perturbations with all harmonics $k$ of the resonant perturbation $(kn)/(km)=\iota$ with $k$ any integer (Huang-Hudson:2022). In a torus, the toroidicity produces magnetic perturbations with poloidal harmonics of strength $\epsilon^{k}$ at $m\pm k$ with $\epsilon$ the inverse aspect ratio of the torus. Any rational number can be produced by the harmonics of $n/m$ together with the harmonics of $m$ produced by toroidicity. The more ideal the perturbation, the more contorted the magnetic surfaces become, which leads to their rapid breaking by an arbitrarily small non-ideal effect, such as resistivity (Boozer:2022). This is discussed in Section 6.
Section 7 gives additional examples of the application if the concepts of magnetic field line chaos, cantori, and turnstiles to toroidal fusion plasmas but also expresses concern about the future of innovation within U. S. fusion program.
@@ 2 Representation of a magnetic field and its field lines
(Boozer:1983) showed that the magnetic field in a toroidal region can always be written as
@@
In general the poloidal flux depends on all three spatial coordinates and time, $\psi_{p}(\psi_{t},\theta,\varphi,t)$. (Boozer:1983) showed the trajectories of the magnetic field lines are given by a Hamiltonian $\psi_{p}(\psi_{t},\theta,\varphi,t)$ in canonical form with time $t$ just a parameter:
@@
The distance between neighboring irrational magnetic surfaces $\Delta(\theta,\varphi)\equiv|\psi_{t}|/|\vec{\nabla}\psi_{t}|$ can have an exponentially large ratio between its maximum and minimun, $\Delta_{max}/\Delta_{min}=e^{2\sigma}$, despite all the surfaces remaining perfect. As discussed in Section 6, when $\sigma\gtrsim 20$ even the small resistivity in fusion plasmas will cause the magnetic surfaces to break and form a region of volume-filling chaotic magnetic field lines. Since $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi$ varies little over a magnetic surface compared to $e^{20}$, places where neighboring surfaces are close together must be places where the field lines in the surfaces are far apart and vice versa (Boozer:2022).
@@ 3 Breakup of magnetic surfaces
An island chain forms at every rational surface that resonates with the perturbation. Realistic magnetic perturbations have many Fourier harmonics. Only in a cylinder with two symmetry directions is a single Fourier harmonic perturbation possible. (Chirikov:1960) noted that when island chains from different rational surfaces overlap the region between the rational surfaces becomes chaotic.
@@ 4 Chaos
Chaos in the sense of exponentiation allows a magnetic field $\vec{B}(\vec{x})$ that has variations only on long spatial scales to have structures in the magnetic field lines on arbitrarily small spatial scales. As illustrated in Figure 3, the exponentiation in separation of neighboring lines causes an exponentially large distortions in any tube of magnetic flux including the tubes formed by neighboring irrational magnetic surfaces. This distortion is of fundamental importance to the physics of magnetic reconnection, Section 6, though generally ignored in the reconnection literature. Neither localized currents nor localized magnetic fields are required. Chaos is observed in curl-free magnetic fields that depend on all three spatial coordinates and are not carefully designed to avoid chaos. Nevertheless, a large exponentiation while field lines lie on perfect magnetic surfaces does not seem possible in a curl-free field.
Chaotic magnetic field lines in which a single field line covers the whole chaotic volume also separate exponentially when followed far enough in $\ell$. But, the existence of cantori with highly collimated turnstiles makes the implication of the first property of chaotic field lines by the second also subtle. As discussed by (MacKay:1984) and in the Introduction, the formation of chaotic regions in which a single line covers a volume involves cantori, which can appear to be just an irrational magnetic surface for an arbitrarily long distance $\ell$ along one of the field lines in the surface. These subtleties are also discussed in Section 5 and must be understood to determine the relevance of magnetohydrodynamic (MHD) simulations to actual plasma behavior.
A subtlety of a field line filling a volume is illustrated by a point made by (Meiss:2015). Although the red area in Figure 2 was formed by a single chaotic trajectory during a million iterations, that trajectory escaped from the red area after approximately a billion iterations. The implication is that the boundary of the red area was a cantorus penetrated by turnstiles that occupy only a tiny fraction of the area of the boundary. (Punjabi:2025) found that outermost confining magnetic surface of a stellarator has a similar subtlety. A region with a width of a few per cent of the $\psi_{t}$ enclosed by the outermost confining surface confines magnetic field lines while they make transits through hundreds of thousands of stellarator periods, but they eventually strike the wall. There is presumably little difference in the behavior of the plasma on these lines than that on lines that have perfect confinement.
@@
When the electron pressure, $P_{e}$, has a gradient along chaotic magnetic field lines, (Boozer:2024:Electric) showed that the electric field required for quasi-neutrality gives $\vec{E}\times\vec{B}$ flows that lead to plasma transport at a Bohm-like rate. A variation in the electric potential, $en\partial\Phi/\partial\ell=\partial P_{e}/\partial\ell$, where $n$ is the electron number density, is required to keep the rapidly moving electrons together with the slowly moving ions. The exponentially large dependence of the separation of neighboring field lines with $\ell$ leads to increasing $\vec{E}\times\vec{B}$ flows until a balance in electron and ion transport is achieved. The required Bohm-like level of transport across the magnetic field lines is important for sweeping in impurities in tokamak disruptions and for transport in non-resonant stellarator divertors.
Velocities can also be chaotic by having chaotic streamlines. In a chaotic flow, each streamline, $\vec{x}(t)$ with $d\vec{x}/dt=\vec{u}(\vec{x},t)$, has neighboring streamlines that increase their separation exponentially with time. A chaotic magnetic field requires a dependence on all three spatial coordinates. A chaotic magnetic field has chaotic magnetic field lines, which are defined at fixed instants in time. However, a chaotic flow requires a dependence on only two spatial coordinates when the flow is time dependent. The failure to appreciate the implications explains the origins of fundamental misconceptions in the theory of magnetic reconnection, Section 6. As discussed in (Boozer:2025:Magnetic), when the flow velocity of magnetic field lines is chaotic, the ideal evolution of a magnetic field in a system with only two spatial coordinates produces an exponential increase of the field strength. When the field depends on all three spatial coordinates, a chaotic flow velocity makes the field lines chaotic while producing only a moderate increase in the magnetic field strength.

### v3

@@ Abstract
The mathematical concepts of chaos, cartori, and turnstiles underlie a number of areas of tokamak and stellarator physics. Nevertheless, few authors have explicitly used these concepts in publications on fusion plasmas. The small number of plasma physicists who have an intuitive understanding of these concepts retards or blocks development in a number of areas: magnetic reconnection, the most important electromagnetic correction to what are called electrostatic micro-instabilities, non-resonant divertors in stellarators, disruptions and damage from runaway electrons in tokamaks. Physicists become interested in new mathematical concepts when they give insights into and solutions to practical problems. The importance of this review is not only in explaining chaos, cartori, and turnstiles as mathematical concepts but also in illustrating their significance through applications.
@@ 1 Introduction
Positions in space are fundamentally defined by Cartesian coordinates, $\vec{x}=x\hat{x}+y\hat{y}+z\hat{z}$. Any three quantities $(\psi_{t},\theta,\varphi)$ can be used as coordinates by defining the position function $\vec{x}(\psi_{t},\theta,\varphi)$ as long as the Jacobian
$\mathcal{J}\equiv\left(\frac{\partial\vec{x}}{\partial\psi_{t}}\times\frac{\partial\vec{x}}{\partial\theta}\right)\cdot\frac{\partial\vec{x}}{\partial\varphi}$ (1)
is neither zero nor infinity in the spatial region in which they are to be used. Defining the position function just means giving $x$, $y$, and $z$ as functions of $(\psi_{t},\theta,\varphi)$. The appendix to (Boozer, 2004) explains how arbitrary coordinate systems, such as $\vec{x}(\psi_{t},\theta,\varphi)$, are used in calculations.
@@
(Boozer, 1983) showed the topological properties of magnetic field lines at each instant in time are given by $\psi_{p}(\psi_{t},\theta,\varphi,t)$. The position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ is needed to plot the field lines in ordinary space with $\partial\vec{x}/\partial t$ giving the velocity of $(\psi_{t},\theta,\varphi)$ through space, which is the field line velocity when field line topology is not changing, which implies $\psi_{p}$ can be taken to be independent of time.
When a particular line $\vec{x}_{0}(\ell)$ remains within a bounded region of space, it has three possibilities as $\ell\rightarrow\infty$: It can close on itself after a distance $\ell_{0}$. It can cover a surface coming arbitrarily close to every point on a surface without ever going through the same point twice. It can cover a volume coming arbitrarily close to every point in a volume of space without ever going through the same point twice. All three types of trajectories are illustrated by the Standard Map, Figure 2. The iteration number of the map represents the distance $\ell$. The periodicity of both coordinates of the Standard Map makes it non-trivial to translate them into the $(\psi_{t},\theta)$ coordinates of the magnetic field line Hamiltonian $\psi_{p}$. (Boozer and Punjabi, 2018) in their Appendix B have obtained an explicit formula for a $\psi_{p}(\psi_{t},\theta,\varphi)$ that represents the qualitative features of stellarator equilibria including the non-resonant divertor that (Strumberger, 1992) found naturally arises at the edge optimized stellarators. Non-resonant divertors are discussed in Section 5.2 as way to build intuitive understanding of the subtleties of cantori and turnstiles.
@@
The behavior of neighboring magnetic field lines, which are infinitesimally separated from an arbitrary line $\vec{x}_{0}(\ell)$, can be studied without approximation by a particularly simple Hamiltonian (Boozer, 2012). The study of neighboring lines, which is reviewed in Appendix A, showed: (1) It is possible for field lines infinitesimally separated from a given line $\vec{x}_{0}(\ell)$, of any of the three types, to have separations that increase or decrease exponentially with $\ell$. There must be an equal flux of neighboring lines that diverge from and approach the line $\vec{x}_{0}(\ell)$ due to $\vec{\nabla}\cdot\vec{B}=0$. When lines that lie on a circle of infinitesimal radius at $\ell=0$ are followed a distance $\ell$, the circle will deform into an ellipse. The field strength times the square of the semi-major axis of the ellipse increases as $e^{\sigma_{max}(\ell)}$, and the field strength times the square semi-minor axis decreases as $e^{-\sigma_{max}(\ell)}$. (2) When the magnetic field depends on only two spatial coordinates, a single field line cannot cover a volume, as in the red area of Figure 2. Three spatial coordinates are required. Two-coordinate magnetic fields, as in axisymmetry, can have X-points (more correctly X-lines) which are specific field lines from which neighboring lines exponentially approach and separate. However, the volume in which lines have at least a certain exponentiation is itself exponentially small. (3) Curl-free magnetic fields can exhibit field lines with an exponential separation throughout a volume. An ideal magnetic evolution does not conserve the exponentiation $\sigma_{max}(\ell)$. In other words, the exponentiation can be a property of the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ with $\psi_{p}$ an unchanging function of its canonical coordinates. The current density required for an ideal evolution to produce an exponentiation $\sigma_{max}$ is proportional to $\sigma_{max}$, Appendix A.3. As will be discussed in Section 6.1.2, the time required for a large scale reconnection to occur is determined by a timescale set by the ideal evolution multiplied by $\sigma_{r}$. The required exponentiation $e^{\sigma_{r}}$ for large scale reconnection is the ratio of timescale for non-ideal diffusive effects, such as resistivity, to a timescale determined by the ideal evolution.
@@
The tendency to construct two dimensional models has led to the prejudice that exponential separation of magnetic field lines throughout a volume is peculiar. Actually any magnetic field that depends on all three spatial coordinates tends to be chaotic, which is illustrated by the design of a curl-free magnetic fields for a stellarators. Obtaining nested magnetic surfaces, as required for a three-dimensional stellarator, rather than having the field lines fill a volume requires careful optimization. Even a small perturbation to an optimized curl-free stellarator can destroy the surfaces. Nevertheless, axisymmetric tokamaks are even more sensitive to external field perturbations due to their amplification by coupling to marginally stable current-driven kinks (Park et al., 2010).
@@
(Boozer, 2004) showed the time derivative of the poloidal flux is the loop voltage,
@@
At the magnetic axis, $\psi_{t}=0$, Equation (2) is a trivial consequence of Faraday’s Law plus Stokes’ Theorem with the loop voltage $V_{\ell}$ the integral of the electric field around the axis. The proof of this equation for an arbitrary magnetic surface is non-trivial but holds as long as the surface enclosing the toroidal flux $\psi_{t}$ exists. On an arbitrary magnetic surface, the loop voltage is an integral along a field line at an instant of time,
@@
As proven in the paragraph below Equation (17), when $\vec{\nabla}V_{\ell}=0$ the topology of the magnetic field lines cannot change, and their evolution is considered ideal. This is in contrast to ideal magnetohydrodynamics, which requires a zero resistivity $\eta$, and that implies $V_{\ell}=0$. As pointed out by (Boozer, 2025a) and discussed in Section 6.2, the distinction between $\vec{\nabla}V_{\ell}=0$ and $V_{\ell}=0$ is important when devising strategies for avoiding disruptions in tokamaks. The condition $\vec{\nabla}V_{\ell}=0$ is consistent with changing the magnetic flux at the axis of a toroidal plasma without changing the magnetic field in the plasma while the condition $V_{\ell}=0$ is not.
@@
Non-resonant perturbations can be taken to change the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ while the poloidal flux can be held fixed using a canonical transformation or what is mathematically equivalent an electric potential. This is shown in the appendix to (Boozer, 2004).
Section 4 discusses the subtle properties of chaotic magnetic field lines. These properties need to be understood before undertaking a study of cantori and turnstiles, which serve as boundaries on regions of chaotic magnetic field lines.
(Chirikov, 1960) gave the approximate condition for the creation of large chaotic volumes: the overlap of the islands from different resonant rational surfaces.
As discussed in Section 5, the actual formation of chaotic regions is more complicated than just the overlap of island chains. (MacKay, Meiss, and Percival, 1984) showed that chaotic regions are formed by cantori and the turnstiles that penetrate them.
@@
The highly collimated flux tubes of turnstiles are responsible for the severe damage that can be caused by runaway electrons following the lines in these tubes to localized positions on the chamber walls (Boozer and Punjabi, 2016). This collimation also responsible for the formation of a natural non-resonant divertor at the edge of a stellarator (Strumberger, 1992) and (Boozer and Punjabi, 2018). Section 5.2 on non-resonant divertors provides an intuitive introduction to cantori and turntiles through an application to a practical problem.
Section 6 is a discussion of applications. The first topic is magnetic reconnection. If the plasma evolution were perfectly ideal, a resonant perturbation would affect only the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$. But, a resonant ideal perturbation forces the magnetic field to have a delta-function current density at the resonant rational surface $\iota=n/m$. This current causes distortions to the neighboring surfaces producing magnetic perturbations with all harmonics $k$ of the resonant perturbation $(kn)/(km)=\iota$ with $k$ any integer (Huang et al., 2022). In a torus, the toroidicity produces magnetic perturbations with poloidal harmonics of strength $\epsilon^{k}$ at $m\pm k$ with $\epsilon$ the inverse aspect ratio of the torus. Any rational number can be produced by the harmonics of $n/m$ together with the harmonics of $m$ produced by toroidicity. The more ideal the perturbation, the more contorted the magnetic surfaces become, which leads to their rapid breaking by an arbitrarily small non-ideal effect, such as resistivity (Boozer, 2022). Section 6 also discusses applications to tokamak disruptions and to stellarators.
Section 7 is an overview of the paper. The Appendix reviews the behavior of magnetic field lines that are infinitesimally separated from an arbitrary given line, $\vec{x}_{0}(\ell)$.
@@ 2 Representation of a magnetic field and its field lines
(Boozer, 1983) showed that the magnetic field in a toroidal region can always be written as
@@
In general the poloidal flux depends on all three spatial coordinates and time, $\psi_{p}(\psi_{t},\theta,\varphi,t)$. (Boozer, 1983) showed the trajectories of the magnetic field lines are given by a Hamiltonian $\psi_{p}(\psi_{t},\theta,\varphi,t)$ in canonical form with time $t$ just a parameter:
@@
The distance between neighboring irrational magnetic surfaces $\Delta(\theta,\varphi)\equiv|\psi_{t}|/|\vec{\nabla}\psi_{t}|$ can have an exponentially large ratio between its maximum and minimun, $\Delta_{max}/\Delta_{min}=e^{2\sigma}$, despite all the surfaces remaining perfect. As discussed in Section 6.1, when $\sigma\gtrsim 20$ even the small resistivity in fusion plasmas will cause the magnetic surfaces to break and form a region of volume-filling chaotic magnetic field lines. Since $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi$ varies little over a magnetic surface compared to $e^{20}$, places where neighboring surfaces are close together must be places where the field lines in the surfaces are far apart and vice versa (Boozer, 2022).
@@ 3 Breakup of magnetic surfaces
An island chain forms at every rational surface that resonates with the perturbation. Realistic magnetic perturbations have many Fourier harmonics. Only in a cylinder with two symmetry directions is a single Fourier harmonic perturbation possible. (Chirikov, 1960) noted that when island chains from different rational surfaces overlap the region between the rational surfaces becomes chaotic.
@@ 4 Chaos
Chaos in the sense of exponentiation allows a magnetic field $\vec{B}(\vec{x})$ that has variations only on long spatial scales to have structures in the magnetic field lines on arbitrarily small spatial scales. As illustrated in Figure 3, the exponentiation in separation of neighboring lines causes an exponentially large distortion in any tube of magnetic flux including the tubes formed by neighboring irrational magnetic surfaces. This distortion is of fundamental importance to the physics of magnetic reconnection, Section 6.1, though generally ignored in the reconnection literature. Neither localized currents nor localized magnetic fields are required. Chaos is observed in curl-free magnetic fields that depend on all three spatial coordinates and are not carefully designed to avoid chaos. Nevertheless, a large exponentiation while field lines lie on perfect magnetic surfaces does not seem possible in a curl-free field.
Chaotic magnetic field lines in which a single field line covers the whole chaotic volume also separate exponentially when followed far enough in $\ell$. But, the existence of cantori with highly collimated turnstiles makes the implication of the first property of chaotic field lines by the second also subtle. As discussed by (MacKay, Meiss, and Percival, 1984) and in the Introduction, the formation of chaotic regions in which a single line covers a volume involves cantori, which can appear to be just an irrational magnetic surface for an arbitrarily long distance $\ell$ along one of the field lines in the surface. These subtleties are also discussed in Section 5 and must be understood to determine the relevance of magnetohydrodynamic (MHD) simulations to actual plasma behavior.
A subtlety of a field line filling a volume is illustrated by a point made by (Meiss, 2015). Although the red area in Figure 2 was formed by a single chaotic trajectory during a million iterations, that trajectory escaped from the red area after approximately a billion iterations. The implication is that the boundary of the red area was a cantorus penetrated by turnstiles that occupy only a tiny fraction of the area of the boundary. (Punjabi and Boozer, 2025) found that outermost confining magnetic surface of a stellarator has a similar subtlety. A region with a width of a few per cent of the $\psi_{t}$ enclosed by the outermost confining surface confines magnetic field lines while they make transits through hundreds of thousands of stellarator periods, but they eventually strike the wall. There is presumably little difference in the behavior of the plasma on these lines than that on lines that have perfect confinement.
@@
When the electron pressure, $p_{e}$, has a gradient along chaotic magnetic field lines, (Boozer, 2024a) showed that the electric field required for quasi-neutrality gives $\vec{E}\times\vec{B}$ flows that lead to plasma transport at a Bohm-like rate. A variation in the electric potential, $en\partial\Phi/\partial\ell=\partial p_{e}/\partial\ell$, where $n$ is the electron number density, is required to keep the rapidly moving electrons together with the slowly moving ions. The exponentially large dependence of the separation of neighboring field lines with $\ell$ leads to increasing $\vec{E}\times\vec{B}$ flows until a balance in electron and ion transport is achieved. The required Bohm-like level of transport across the magnetic field lines is important for sweeping in impurities in tokamak disruptions and for transport in non-resonant stellarator divertors.
The velocity of a flow can also be chaotic by having chaotic streamlines. In a chaotic flow, each streamline, $\vec{x}(t)$ with $d\vec{x}/dt=\vec{u}(\vec{x},t)$, has neighboring streamlines that increase their separation exponentially with time. A chaotic magnetic field requires a dependence on all three spatial coordinates. A chaotic magnetic field has chaotic magnetic field lines, which are defined at fixed instants in time. However, a chaotic flow requires a dependence on only two spatial coordinates when the flow is time dependent. The failure to appreciate the implications explains the origins of fundamental misconceptions in the theory of magnetic reconnection, Section 6.1. As discussed in (Boozer, 2025c), when the flow velocity of magnetic field lines is chaotic, the ideal evolution of a magnetic field in a system with only two spatial coordinates produces an exponential increase of the field strength. When the field depends on all three spatial coordinates, a chaotic flow velocity makes the field lines chaotic while producing only a moderate increase in the magnetic field strength.